OpenStudy (idku):

Runaway ski slides down a 260-m-long slope inclined at 36º with the horizontal. If the initial speed is 12.0 m/s, how long does it take the ski to reach the bottom of the incline if the coefficient of kinetic friction between the ski and snow is 0.100?

OpenStudy (idku):

Given: \(\color{#000000}{ \displaystyle d=260 ~{\sf m} }\) \(\color{#000000}{ \displaystyle \theta=36^{\circ} }\) \(\color{#000000}{ \displaystyle v_0=12.0 ~{\sf m/s} }\) \(\color{#000000}{ \displaystyle \mu_k=0.100 }\) Looked for: \(\color{#000000}{ \displaystyle t=?~{\sf s} }\) (The time to reach the bottom of this 260-m slope )